OpenStudy (anonymous):

is this RIGHT..?? Arrange the steps in the correct sequence to find the inverse function of f(x)=3x-15/2

OpenStudy (anonymous):

here

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

Follow these steps and you'll get it :)

ganeshie8 (ganeshie8):

\[\begin{align} f(x) = y &= \dfrac{3x-15}{2}\\ &\downarrow\\ \end{align}\]

ganeshie8 (ganeshie8):

step1 looks good as you have started with the given function

Miracrown (miracrown):

Wait...you started with (3x - 15)/2 but the 2nd line has 3x + 15 = 2y?

ganeshie8 (ganeshie8):

yeah second step should be replaced

OpenStudy (anonymous):

i know @Miracrown is weird

Miracrown (miracrown):

Oh, you have other choices in the tiles

OpenStudy (anonymous):

Look at what I said after replacing f(x) with y, second step is wrong. Switch x's and y's, so put x where y is and x where y is.

ganeshie8 (ganeshie8):

in step2 you swap x and y

Miracrown (miracrown):

Yes, let me look at it again...

OpenStudy (anonymous):

so which one is that?

OpenStudy (anonymous):

Look at your tiles

Miracrown (miracrown):

The first step should be to switch x and y That's what an inverse means So I'm thinking x = (3y - 15)/2 is first

ganeshie8 (ganeshie8):

\[\begin{align} f(x) = \color{green}{y} &= \dfrac{3\color{Red}{x}-15}{2}\\ &\downarrow\\ \color{Red}{x} &= \dfrac{3\color{green}{y}-15}{2}\\ \end{align}\]

ganeshie8 (ganeshie8):

i would like to put a box around steps hmm but latex is not working

OpenStudy (anonymous):

draw it out

ganeshie8 (ganeshie8):

i want to draw it out in latex

OpenStudy (anonymous):

got that :)

OpenStudy (anonymous):

ganeshie xD

Miracrown (miracrown):

http://prntscr.com/51u94f If you are allowed to change that first tile, then the first step should be to say that f(x) = y And then switch x and y

OpenStudy (anonymous):

Your third step is wrong man

OpenStudy (anonymous):

yep i did it :)..so what abot step 3,4,5

OpenStudy (anonymous):

You have to solve for y not x

OpenStudy (anonymous):

i knew it -_-

OpenStudy (anonymous):

So lets put it in order here

OpenStudy (anonymous):

Step 1. \[f(x) = y = \frac{ 3x-15 }{ 2 }\] Step 2. \[x = \frac{ 3y-15 }{ 2 }\] Now lets solve for y. Step 3. \[2x = 3y-15\]

OpenStudy (anonymous):

Now what would be next?

OpenStudy (anonymous):

add them? or subtract?

OpenStudy (anonymous):

Add the 15 on both sides, what do you get?

OpenStudy (anonymous):

go away dan

Miracrown (miracrown):

I probably wouldn't look at the choices yet...I would solve the whole thing for the inverse first, then look at the choices. The boxes get too confusing since they threw in "wrong" steps to throw you off

OpenStudy (dan815):

inverse is u wanna do the steps backwards such that u get back what u started with

OpenStudy (anonymous):

2x+15=3y

OpenStudy (anonymous):

Yeah good job @cmvbubu

OpenStudy (anonymous):

What's after that?

OpenStudy (anonymous):

this is crazy but i apreciate everybodys help in here !:D

OpenStudy (dan815):

if y= f(x) what shud x be then the function that is giving back the x now is called the inverse function x=f'-1(y) but we say hey y is the unknown and x is the input so y=f^-1(x)

OpenStudy (anonymous):

Step 4. \[3y = 2x+15\]

OpenStudy (anonymous):

so u end up with f^-1(x)=y=2x+15/3

OpenStudy (dan815):

now given some input u get to know y, which is the value u must have put into the forward function for x to get back the x u put in when is the y from that one

OpenStudy (anonymous):

Yay good job @cmvbubu :)

OpenStudy (dan815):

now given some input u get to know y, which is the value u must have put into the forward function for x to get back the x u put in which* is the y from that one

OpenStudy (dan815):

perfect explaination i know ty danke

OpenStudy (dan815):

"dan"ke

OpenStudy (anonymous):

dan you were a fat baby

OpenStudy (anonymous):

wait i think all babies are fat

OpenStudy (dan815):

This is abb0t not me

OpenStudy (anonymous):

Oh makes sense

OpenStudy (anonymous):

i change step 4?

OpenStudy (anonymous):

i just did

OpenStudy (anonymous):

Your step 3 was wrong

OpenStudy (anonymous):

step 4 and 5 were right

OpenStudy (anonymous):

SO basically 2 and 3 were wrong for you

OpenStudy (anonymous):

so it should look like this

OpenStudy (anonymous):

So again, To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x) if you follow these steps all of these questions are pretty much the same.

OpenStudy (anonymous):

Your step 4 is wrong now

OpenStudy (anonymous):

waa lol ok hold on....

OpenStudy (anonymous):

It should be 2x+15=3y don't know why you did that.

OpenStudy (anonymous):

oh yeah i was confuse i didnt look at the variables

OpenStudy (anonymous):

details suck

Miracrown (miracrown):

No, not really.

OpenStudy (anonymous):

It can get confusing since you're dealing with inverses but after a couple of problems you'll realize it's usually the same, as far as you need to know for now.

OpenStudy (anonymous):

DETAILS are important, that's how you get part marks ;)

OpenStudy (dan815):

details r for n00bs

Miracrown (miracrown):

Yeah, n00bs like u - dan

OpenStudy (anonymous):

so this is the right one?

OpenStudy (anonymous):

Looks good

OpenStudy (dan815):

the negative of dan is a n00b so the positive is so pro

Miracrown (miracrown):

Pfft.

OpenStudy (anonymous):

Cage match, dan vs Mira

OpenStudy (dan815):

im not entering a cage with that witch on halloween

OpenStudy (anonymous):

so last question why u so smart?? dont tell me cause you are BATMAN!!

OpenStudy (anonymous):

loool

Miracrown (miracrown):

Dan: 0 Mira: 1

OpenStudy (dan815):

Mira: i

OpenStudy (anonymous):

Me smart? Omg I'm so flattered, I'm an idiot these other people are geniuses :p

OpenStudy (anonymous):

Thanks everybody here new Fan :)

Miracrown (miracrown):

No one is smarter than me ;)

OpenStudy (dan815):

on the imaginary axis!

Miracrown (miracrown):

Shhhh

OpenStudy (dan815):

bahahah

Miracrown (miracrown):

I bet ganeshie is writing a lecture for us

OpenStudy (anonymous):

http://mathworld.wolfram.com/Witchoowletnesi.html Need to mention this before halloween ends!

OpenStudy (anonymous):

Least ganeshie got the box's to work

Miracrown (miracrown):

It was ended 18 hours ago.

OpenStudy (anonymous):

12 minutes left here :)

Miracrown (miracrown):

Make them mins evil for everyone else.

OpenStudy (anonymous):

why when u post a comment is says ''an hour ago'' when i know it was just a second ago?lol

OpenStudy (anonymous):

The time stamp is messed up, some posts say yesterday when it was 5 minutes ago

Miracrown (miracrown):

Welcome to the OS ''Glitch'' Club.

OpenStudy (anonymous):

^

OpenStudy (anonymous):

jhajah ^ ok wow thats messup

Miracrown (miracrown):

Indeed.